OpenStudy (anonymous):

Base of a triangle is 2 1/3 and height 3/5 yard. The area of a triangle equals half the product of its base and height. = _____yd^2

OpenStudy (lgbasallote):

\[\Large A = \frac{bh}{2} \implies \frac{ 2 \frac 13 \times \frac 35}{2} \implies \frac{ \frac 73 \times \frac 35}{2} \implies \frac{ \frac 75}{2} \implies \frac{7}{10}\] do you get that?

OpenStudy (lgbasallote):

feel free to ask anything about my solution